OpenStudy (anonymous):

The line that passes through (-3, 5) and (-2, -6)

OpenStudy (anonymous):

slope = (5-(-6)) / (-3-(-2)) =11/-1 =-11 (y-5)/(x-(-3)) =-11 y-5=(-11)(x+3) y-5=-11x-33 11x-y+28=0

OpenStudy (anonymous):

After getting it into Point-Slope form, y - y1 = m(x - x1),you need to continue solving for y to get it into Slope-Intercept form. y - 5 = -11(x + 3) y - 5 = -11x - 33 Add 5 to both sides. y = -11x - 28 ^That is the equation for your line.